OpenStudy (anonymous):

3sin^2 theta-sin theta-2=0

OpenStudy (campbell_st):

this is basically a quadratic equation that can be factored \[(3\sin(\theta)+ 2 )(\sin(\theta -1) = 0\] so then you need to solve 3sin(@) + 2 = 0 and sin(@) -1 = 0 hope it helps